A large mining company was provided with a 3 cu. meters of compressed air tank. Air pressure in the tank drops from 700 kPa to 180 kPa while the temperature remains unchanged at 28 0C. What percentage has the mass of air in the tank been reduced.


1
Expert's answer
2021-10-19T15:13:36-0400

Ideal gas law

"pV=nRT \\

pV= \frac{m}{M}RT \\

m = \frac{pVM}{RT}"

Initial mass:

"m_i = \frac{p_iVM}{RT}"

Final mass:

"m_f = \frac{p_fVM}{RT} \\

Δm=m_i -m_f = (p_i- p_f) \frac{VM}{RT}"

Percantage mass reduced:

"= \frac{Δm}{m_i} \times 100 = \frac{(p_i-p_f) \frac{VM}{RT}}{p_i \frac{VM}{RT}} \times 100 \\

= \frac{p_i -p_f}{p_i} \times 100 \\

= \frac{700-180}{700} \times 100 \\

= 74.2 \; \%"